# AirSketch: Generative Motion to Sketch

### v1

@@ Abstract
Illustration is a fundamental mode of human expression and communication. Certain types of motion that accompany speech can provide this illustrative mode of communication. While Augmented and Virtual Reality technologies (AR/VR) have introduced tools for producing drawings with hand motions (air drawing), they typically require costly hardware and additional digital markers, thereby limiting their accessibility and portability. Furthermore, air drawing demands considerable skill to achieve aesthetic results. To address these challenges, we introduce the concept of AirSketch, aimed at generating faithful and visually coherent sketches directly from hand motions, eliminating the need for complicated headsets or markers. We devise a simple augmentation-based self-supervised training procedure, enabling a controllable image diffusion model to learn to translate from highly noisy hand tracking images to clean, aesthetically pleasing sketches, while preserving the essential visual cues from the original tracking data. We present two air drawing datasets to study this problem. Our findings demonstrate that beyond producing photo-realistic images from precise spatial inputs, controllable image diffusion can effectively produce a refined, clear sketch from a noisy input. Our work serves as an initial step towards marker-less air drawing and reveals distinct applications of controllable diffusion models to AirSketch and AR/VR in general.
@@ 1 Introduction
Our experiments show that with our augmentation-based training, controllable image DMs are able to recognize and interpret correct visual cues from noisy tracking images, some of which even appear to be nearly unrecognizable to the human eye, and generate sketches faithfully and aesthetically, while being robust to unseen objects. Moreover, we show that through simple augmentations, the model is capable of sketch-completion and text-instructed stroke styling. Finally, we conduct ablations to investigate 1) the effects of different augmentations, 2) the contribution of text prompts to sketch generation and 3) the effect of different level of input ‘chaos’ on the quality of resulting generations.
@@ 2.2 Sketch Applications
Most existing sketch generation methods adopt the vector representation [22, 55, 37], and view sketch generation as an auto-regressive problem on coordinate sequences. These sequences are typically represented by lists of tuple $(\delta x,\delta y,\mathbf{p})$, where $\delta x$ and $\delta y$ represent the offset distances of x and y from the previous point, and $\mathbf{p}$ is a one-hot vector indicating the state of the pen. Sketch-RNN [22] uses a Variational Autoencoder (VAE) with a bi-directional RNN as the encoder and an auto-regressive RNN as the decoder. Sketch-Bert [37] follows the BERT [14] model design and training regime. Instead of auto-regressive sketch generation, SketchGAN [38] takes in a rasterized sketch image and uses a cascade GAN to iteratively complete the sketch.
@@ 4.3 Sketch Augmentations
Table 1: Results on the similarity between generated and ground-truth sketches. “Tracking” refers to hand tracking images, and “Gen.” refers to generated images. “w/ Aug.” refers to whether sketch augmentations have been applied. “CLIP I2I/I2T” refers to CLIP Image-to-Image/Image-to-Text similarity.
@@
5 Experiments
Datasets and Implementation Details.
Figure 5: Inference results on unseen categories. For the left group, the categories are: angel, cat, cow, sun; for the right group, the classes are: cow, candle, car, snail (top to bottom).
Evaluation Metrics.
@@ 5.1 Results and Analysis
In Figure 1 and Figure 6 we can clearly observe the ControlNet trained with augmentations successfully identifies the visual cues from the noisy tracking image, removes artifacts, and generates the clean sketches, while being aesthetic and semantically coherent. Unsurprisingly, ControlNet trained without augmentations fails to make any improvement from the tracking.
@@ 6 Conclusions
In this work we provide a stepping stone for the proposed task by creating a framework that establishes a correspondence between hand motions and clean, pleasing, and coherent illustrations, and providing datasets to evaluate these tasks. However, this work does not explore the possibility of using hand gestures to create complex, full color images that image diffusers are known for. It also assumes that the desired output sketch is simple and often cartoon-like, as a majority of the Quick, Draw! sketches are drawn with few lines and simple shapes. These are interesting future research directions.

### v3

@@ Abstract
Illustration is a fundamental mode of human expression and communication. Certain types of motion that accompany speech can provide this illustrative mode of communication. While Augmented and Virtual Reality technologies (AR/VR) have introduced tools for producing drawings with hand motions (air drawing), they typically require costly hardware and additional digital markers, thereby limiting their accessibility and portability. Furthermore, air drawing demands considerable skill to achieve aesthetic results. To address these challenges, we introduce the concept of AirSketch, aimed at generating faithful and visually coherent sketches directly from hand motions, eliminating the need for complicated headsets or markers. We devise a simple augmentation-based self-supervised training procedure, enabling a controllable image diffusion model to learn to translate from highly noisy hand tracking images to clean, aesthetically pleasing sketches, while preserving the essential visual cues from the original tracking data. We present two air drawing datasets to study this problem. Our findings demonstrate that beyond producing photo-realistic images from precise spatial inputs, controllable image diffusion can effectively produce a refined, clear sketch from a noisy input. Our work serves as an initial step towards marker-less air drawing and reveals distinct applications of controllable diffusion models to AirSketch and AR/VR in general. Code and dataset are available under: https://github.com/hxgr4ce/DoodleFusion.
@@ 1 Introduction
Our experiments show that with our augmentation-based training, controllable image DMs are able to recognize and interpret correct visual cues from noisy tracking images, some of which even appear to be nearly unrecognizable to the human eye, and generate sketches faithfully and aesthetically, while being robust to unseen objects. Moreover, we show that through simple augmentations, the model is capable of sketch-completion and text-instructed stroke styling. Finally, we conduct ablations to investigate 1) the effects of different augmentations, 2) the contribution of text prompts to sketch generation and 3) the effect of different levels of input ‘chaos’ on the quality of resulting generations.
@@ 2.2 Sketch Applications
Most existing sketch generation methods adopt the vector representation [22, 55, 37], and view sketch generation as an auto-regressive problem on coordinate sequences. These sequences are typically represented by lists of tuple $(\delta x,\delta y,\mathbf{p})$, where $\delta x$ and $\delta y$ represent the offset distances of x and y from the previous point, and $\mathbf{p}$ is a one-hot vector indicating the state of the pen. Sketch-RNN [22] uses a Variational Autoencoder (VAE) with a bi-directional RNN as the encoder and an auto-regressive RNN as the decoder. Sketch-Bert [37] follows the BERT [14] model design and training regime. Instead of auto-regressive sketch generation, SketchGAN [38] takes in a rasterized sketch image and uses a cascade GAN to iteratively complete the sketch. Sketch-Pix2Seq [sketchpix2seq] adds a vision encoder on top of Sketch-RNN and thus allow the model to take in an image input and reconstruct the sketch using coordinate sequences.
@@ 4.3 Sketch Augmentations
5 Experiments
Datasets and Implementation Details.
Evaluation Metrics.
Table 1: Results on the similarity between generated and ground-truth sketches from Quick, Draw! dataset. “Tracking” refers to hand tracking images, and “Gen.” refers to generated images. “w/ Aug.” refers to whether sketch augmentations have been applied. “CLIP I2I/I2T” refers to CLIP Image-to-Image/Image-to-Text similarity.
@@ 5.1 Results and Analysis
In Figure 1 we can clearly observe the ControlNet trained with augmentations successfully identifies the visual cues from the noisy tracking image, removes artifacts, and generates the clean sketches, while being aesthetic and semantically coherent. Unsurprisingly, ControlNet trained without augmentations fails to make any improvement from the tracking. In Figure 5 we show additional results where the model is trained on TUBerlin [16] dataset.
Figure 5: Generations on TUBerlin dataset.
@@
Table 2: Quantitative comparison between Sketch-Pix2Seq (P2S.) and ours on a subset of 10 classes.
SSIM ($\uparrow$) CD ($\downarrow$) I2I ($\uparrow$) I2T ($\uparrow$)
Tracking 0.5 32.36 0.76 0.21
P2S. 0.58 30.19 0.82 0.26
Ours 0.63 25.45 0.83 0.29
Figure 6: A comparison between generations from Sketch-Pix2Seq (top) and ours (bottom).
In Table 6 and Figure 6 we show comparison between ours and Sketch-Pix2Seq [sketchpix2seq] (P2S). Since P2S performs best when the model is trained on only one category, we randomly pick 10 categories to train 10 P2S models, and compare with ours using the selected categories. For each P2S model, following the original work, we first train 60M steps for reconstructing the exact same input sketch, and subsequently train 40M steps for constructing the clean sketch given the noisy tracking image. In Table 6 our SSIM and CD are noticeably higher than P2S, suggesting better faithfulness. This is validated in Figure 6, where we can observe while P2S is able to generate semantically correct sketches, it fails to follow the input tracking faithfully. Moreover, as P2S requires to train a separate model on each individual category and has no generalizability, it is hard to adapt to real-world usage.
Figure 7: A comparison of visualization of ControlNet hidden states throughout denoising process from baseline approach without augmentation (top) and with augmentations (bottom).
In Figure 7 we show visualizations of the ControlNet hidden states across denoising steps for both the baseline model, trained with the original ControlNet recipe (top row), and our model with noise-augmented training (bottom row). The baseline ControlNet hidden states closely approximate the input edge structure early in the denoising process but fail to converge toward a clean sketch representation, resulting in largely static visualizations. In contrast, the hidden states from our noise-augmented training progressively reveal a coherent and accurate sketch outline throughout the denoising process.
Figure 8: Generations on unseen categories from Quick, Draw! dataset.
@@ 6 Conclusions
We provide a stepping stone for the proposed task by creating a framework that establishes a correspondence between hand motions and clean, pleasing, and coherent illustrations, and providing datasets to evaluate these tasks. However, this work does not explore the possibility of using hand gestures to create complex, full color images that image diffusers are known for. It also assumes that the desired output sketch is simple and often cartoon-like, as a majority of the Quick, Draw! sketches are drawn with few lines and simple shapes.
Societal Impact.
Generative models for media are generally prone to misuse which can lead to the spread of inappropriate and harmful content. While this work is based in the sketch domain, there could still be adverse adoption for generating mocking content and bullying, especially by younger users.